### v2

# VisualQuality-R1: Reasoning-Induced Image Quality Assessment via Reinforcement Learning to Rank

Abstract
DeepSeek-R1 has demonstrated remarkable effectiveness in incentivizing reasoning and generalization capabilities of large language models (LLMs) through reinforcement learning. Nevertheless, the potential of reasoning-induced computation has not been thoroughly explored in the context of image quality assessment (IQA), a task depending critically on visual reasoning. In this paper, we introduce VisualQuality-R1, a reasoning-induced no-reference IQA (NR-IQA) model, and we train it with reinforcement learning to rank, a learning algorithm tailored to the intrinsically relative nature of visual quality. Specifically, for a pair of images, we employ group relative policy optimization to generate multiple quality scores for each image. These estimates are used to compute comparative probabilities of one image having higher quality than the other under the Thurstone model. Rewards for each quality estimate are defined using continuous fidelity measures rather than discretized binary labels. Extensive experiments show that the proposed VisualQuality-R1 consistently outperforms discriminative deep learning-based NR-IQA models as well as a recent reasoning-induced quality regression method. Moreover, VisualQuality-R1 is capable of generating contextually rich, human-aligned quality descriptions, and supports multi-dataset training without requiring perceptual scale realignment. These features make VisualQuality-R1 especially well-suited for reliably measuring progress in a wide range of image processing tasks like super-resolution and image generation.
Figure 1: VisualQuality-R1 excels at image quality scoring, while generating contextually rich, human-aligned quality descriptions.
1 Introduction
Image quality assessment (IQA) aims to quantify the visual quality of digital images consistent with human perceptual judgments. Commonly, IQA models are classified into full-reference (FR) and no-reference (NR) approaches wang2004image , depending on the availability of pristine-quality reference images. In this paper, we focus on NR-IQA due to its practical relevance in real-world scenarios where reference images are unavailable. Over the decades, NR-IQA has evolved from knowledge-driven mittal2012no ; ghadiyaram2017perceptual to data-driven approaches ma2017end ; ke2021musiq ; yang2022maniqa , and shifted from regression-based to ranking-based zhang2021uncertainty ; zhang2023blind techniques. Nevertheless, achieving strong model generalization (e.g., generalization to unseen image distortions) remains a significant, unresolved challenge, driving recent research toward multi-dataset training chen2025toward , active fine-tuning wang2021troubleshooting , and continual model adaptation zhang2022continual .
The rapid advancement of vision-language models (VLMs) offers promising avenues for enhancing NR-IQA generalization by contextualizing it into broader vision tasks wu2024comprehensive . VLMs can effectively integrate multi-modal information, enabling understanding of both low-level image distortions (e.g., noise and blur) and high-level perceptual attributes (e.g., aesthetics and content semantics). This multi-modal semantic contextualization allows VLMs to articulate nuanced quality descriptions with stronger generalization.
However, current NR-IQA methods mainly leverage VLMs through supervised fine-tuning (SFT), which face several critical limitations wu2024qinstruct ; you2024depicting . First, constructing informative quality descriptions demands extensive human effort, rendering the annotation process labor-intensive and expensive [^1] Utilizing state-of-the-art proprietary VLMs such as GPT-4o hurst2024gpt for automated annotation suffers from similar scalability challenges due to high computational costs and financial burdens.. Second, models trained via SFT often overfit to the biases and idiosyncrasies present in training data, and may unintentionally encounter catastrophic forgetting of acquired knowledge during pre-training. Third, SFT typically yields overly rigid and templated outputs (see Fig. 1) that may be less useful.
Reinforcement learning (RL) has recently emerged as a powerful alternative, enhancing the reasoning capabilities of LLMs, while aligning their responses with human preferences rafailov2023direct ; guo2025deepseek . In particular, DeepSeek-R1 guo2025deepseek demonstrates the effectiveness of RL in promoting generalization by encouraging automated exploration of plausible reasoning paths and employing rule-based rewards to prevent reward hacking skalse2022defining . However, a direct adaptation of RL techniques to NR-IQA, as exemplified by the recent Q-Insight model li2025qinsight , has been limited by its reliance on dataset-specific reward design and additional distortion-type classification. These constraints stem from its treatment of visual quality as an absolute perceptual quantity, thereby framing NR-IQA naïvely as a regression task.
In this paper, we introduce VisualQuality-R1, a reasoning-induced NR-IQA model, and we train it via reinforcement learning to rank (RL2R), a learning algorithm explicitly designed to capture the inherently relative nature of visual quality. Specifically, we employ group relative policy optimization (GRPO) shao2024deepseekmath to derive multiple quality scores for each image in a pair. We then compute comparative probabilities between images using the Thurstone model thurstone1927law by assessing the difference between the mean quality score of one image and individual quality scores of another, normalized by their sample variances. Unlike previous methods, we define the reward function using the continuous fidelity measure tsai2007frank , which provides precise guidance to facilitate quality ranking. Extensive experiments confirm that VisualQuality-R1 effectively assesses visual quality across a diverse range of distortion scenarios, outperforming discriminative deep learning-based NR-IQA models as well as a recent reasoning-induced quality regression method li2025qinsight . Moreover, VisualQuality-R1 generates contextually rich, human-aligned quality descriptions (see Fig. 1), which can be leveraged to provide targeted feedback for downstream image processing algorithms, and support fine-grained quality control in digital photography pipelines. Additionally, we demonstrate that VisualQuality-R1 remains effective across multi-dataset training scenarios without requiring perceptual scale realignment.
2 Related Work
This section provides a structured review of related NR-IQA models, emphasizing recent advancements, particularly those leveraging VLMs.
Regression-based Models
NR-IQA models primarily employed regression-based approaches, wherein image quality was treated as an absolute perceptual quantity directly estimated from extracted “quality-aware” features. Initially, features were handcrafted based on natural scene statistics mittal2012no ; mittal2012making , degradation-specific characteristics wang2000blind ; wang2003local ; liu2009no , and perceptual models inspired by the human visual system wang2011reduced . Nonetheless, these methods were limited by the representational capacity of handcrafted features. Later, deep learning-based regression models emerged as the dominant paradigm, using end-to-end trainable neural networks to directly predict quality scores (or, in some cases, quality distributions) kang2014convolutional ; ma2017end ; bosse2017deep ; talebi2018nima ; yang2022maniqa ; wu2023assessor360 . These models typically utilize standard regression losses such as the mean squared error and mean absolute error, or statistical distances such as the earth mover’s distance talebi2018nima and Kullback–Leibler (KL) divergence you2025teaching . Regression-based models often struggle with generalization issues, and require labor-intensive perceptual scale realignment mikhailiuk2021consolidated when training on multiple IQA datasets.
Ranking-based Models
To address these shortcomings, ranking-based NR-IQA models were introduced, modeling visual quality as an intrinsically relative perceptual quantity. Gao et al. gao2015learning pioneered the concept of quality ranking in NR-IQA, although their initial implementation relied on predefined anchor images and was not end-to-end optimized. Ma et al. ma2017dipiq adapted RankNet burges2005learning to NR-IQA by training (though not fully end-to-end) on quality-discriminable image pairs. Their subsequent work established the first end-to-end ranking-based NR-IQA method grounded in the Thurstone model thurstone1927law . Nevertheless, their approach suffers from scaling ambiguity during variance estimation. Zhang et al. zhang2021uncertainty incorporated a hinge loss to regularize variance estimation, yet the scaling ambiguity persisted. Their study also demonstrated the superiority of the fidelity loss tsai2007frank over the conventional cross-entropy loss in ranking-based NR-IQA. Subsequent research has adopted a simpler approach by fixing the variance parameter to one (corresponding to the Thurstone Case V model), facilitating active fine-tuning of NR-IQA models on challenging examples wang2021troubleshooting ; wang2021active and allowing for continual adaptation to novel distortion scenarios zhang2022continual . Other losses that enable quality ranking include the margin ranking loss liu2017rankiqa , differentiable approximations of Spearman’s rank correlation coefficient (SRCC) blondel2020fast , Pearson linear correlation coefficient (PLCC) xu2024boosting , and statistical distances between permutation probabilities cao2007learning ; stern1990models ; irurozki2019mallows .
VLM-based Models
The integration of VLMs into NR-IQA has recently gained traction, particularly due to their proficiency in capturing contextual semantics through multi-modal representation learning. Early attempts include multitask adaptation of CLIP zhang2023blind , as well as SFT-based methods like Q-Align wu2024qalign , Compare2Score zhu2024adaptive , DepictQA you2024depicting , and DeQA-Score you2025teaching , which trained VLMs to generate either quality scores, distributions, or descriptions. Closest to ours, Q-Insight li2025qinsight explored reasoning-induced quality regression through RL. However, Q-Insight struggles with the dataset-specific reward calibration, added complexity of auxiliary distortion-type classification, and generalization to novel distortion scenarios. In contrast, our VisualQuality-R1 redefines the use of VLMs in NR-IQA by shifting from absolute regression to relative ranking, leading to enhanced generalization across distortion scenarios with better quality justifications.
Figure 2: System diagram of the proposed VisualQuality-R1 trained via RL2R. Given an image pair $(x_{i},x_{j})$ with a shared text prompt $c$, VisualQuality-R1 generates $K$ responses. Following GRPO shao2024deepseekmath , each response includes a detailed reasoning process and a predicted quality score. To assess relative visual quality, we calculate the asymmetric comparative probability that image $x_{i}$ is perceived better than $x_{j}$ under the Thurstone model thurstone1927law . This involves subtracting the mean predicted score of $x_{j}$ from the $k$-th score of $x_{i}$, standardized by their sample variances. A fidelity reward is derived from human preference, providing continuous supervisory signals for policy optimization.
3 Reasoning-Induced NR-IQA
To harness both the powerful reasoning-inducing capabilities of RL and the intrinsically relative nature of visual quality, we propose an NR-IQA model—VisualQuality-R1—and an RL2R method of training it that seamlessly integrates the Thurstone model within GRPO. Fig. 2 shows the system diagram of VisualQuality-R1.
3.1 VisualQuality-R1 via RL2R
Given a text prompt $c$ and an image $x$, our goal is to fine-tune a pre-trained VLM, with policy $\pi_{\theta}(\cdot|c,x)$, to produce a scalar quality score in the range of $[1,5]$, following a step-by-step reasoning process, encapsulated within specially designated tags for explicit instruction and enhanced interpretability. The complete structured text prompt is provided in Table 1.
More specifically, for a training batch of images $\{x_{1},x_{2},\ldots,x_{B}\}$, where $B$ is the minibatch size, we apply GRPO to generate $K$ quality predictions for $x_{i}$, $q(x_{i})=[q_{1}(x_{i}),q_{2}(x_{i}),\ldots,q_{K}(x_{i})]^{\intercal}$. This output naturally encodes predictive uncertainty, which is crucial for making reliable relative quality ranking. Under the Thurstone model thurstone1927law , the visual quality of an image is assumed to follow a Gaussian distribution. Thus, we compute the asymmetric comparative probability for each of the $B\times(B-1)$ ordered image pairs by subtracting the mean quality score of $x_{j}$ from the $k$-th quality score of $x_{i}$, standardized by their sample variances:
${p}_{k}(x_{i},x_{j})=\Phi\left(\frac{q_{k}(x_{i})-\mu(q(x_{j}))}{\sqrt{\sigma^{2}(q(x_{i}))+\sigma^{2}(q(x_{j}))+\gamma}}\right),\quad\textrm{for }i\neq j,$ (1)
where $\Phi(\cdot)$ is the standard Gaussian cumulative distribution function. $\mu(q(x_{j}))$ and $\sigma^{2}(q(x_{j}))$ represent the mean and variance of the quality predictions for $x_{j}$, respectively. $\gamma$ is a small positive constant to avoid any potential division by zero. Compared to previous ranking-based NR-IQA models that fix the variance parameter in Eq. (1) to one, we explicitly leverage sample variances derived from GRPO. This gives us an opportunity to dynamically accommodate predictive uncertainty for different images. Meanwhile, using the sample mean for quality comparison stabilizes the asymmetric probability estimate and the subsequent reward calculation by appropriately penalizing outlier predictions.
The true preference $p(x_{i},x_{j})$ is derived from human mean opinion scores (MOSs):
$p(x,y)=\begin{cases}1&\text{if }\mathrm{MOS}(x)>\mathrm{MOS}(y)\\ 0.5&\text{if }\mathrm{MOS}(x)=\mathrm{MOS}(y)\\ 0&\text{otherwise}\end{cases}.$ (2)
Table 1: Structured text prompt used in VisualQuality-R1.
You are doing the image quality assessment task. Here is the question:
What is your overall rating on the quality of this picture? The rating should be a float between 1 and 5, rounded to two decimal places, with 1 representing very poor quality and 5 representing excellent quality. First output the thinking process in <think> </think> tags and then output the final answer with only one score in <answer> </answer> tags.
An important aspect of our RL2R algorithm is that we define the reward function $r_{k}(x_{i})$ for each quality estimate $q_{k}(x_{i})$ as the fidelity measure tsai2007frank —a continuous analogue of the discretized binary reward guo2025deepseek ; li2025qinsight , averaged across all B1B-1 image pairs:
$r_{k}(x_{i})=\frac{1}{B-1}\sum_{j\neq i}\left(\sqrt{p(x_{i},x_{j})p_{k}(x_{i},x_{j})}+\sqrt{(1-p(x_{i},x_{j}))(1-p_{k}(x_{i},x_{j}))}\right).$ (3)
This continuous reward feedback provides precise guidance during RL2R by capturing subtle variations in quality ranking, thus improving generalization across diverse distortion scenarios. We collect $K$ fidelity rewards for $x_{i}$ into the vector $r(x_{i})=[r_{1}(x_{i}),r_{2}(x_{i}),\ldots,r_{K}(x_{i})]^{\intercal}$, and compute the relative advantage $a_{k}(x_{i})$ by standardizing rewards within group:
$a_{k}(x_{i})=\frac{r_{k}(x_{i})-\mu(r(x_{i}))}{\sigma(r(x_{i}))}.$ (4)
The final policy update of $\pi_{\theta}(\cdot|c,x_{i})$ is guided by the regularized objective in GRPO:
$\ell(\theta)=\frac{1}{BK}\sum_{i=1}^{B}\sum_{k=1}^{K}\Bigg($ $\min\left(\frac{\pi_{\theta}(o_{k}|c,x_{i})}{\pi_{\theta_{\text{old}}}(o_{k}|c,x_{i})}a_{k}(x_{i}),\text{clip}\left(\frac{\pi_{\theta}(o_{k}|c,x_{i})}{\pi_{\theta_{\text{old}}}(o_{k}|c,x_{i})},1-\epsilon,1+\epsilon\right)a_{k}(x_{i})\right)$ (5)
$-\beta\,{D}_{\text{KL}}\left(\pi_{\theta}(o_{k}|c,x_{i})\|\pi_{\text{ref}}(o_{k}|c,x_{i})\right)\Bigg).$
Here, $\pi_{\theta_{\text{ref}}}(\cdot|c,x_{i})$ denotes the stable reference policy obtained after VLM pre-training, and $\pi_{\theta_{\text{old}}}(\cdot|c,x_{i})$ is the policy from the previous RL2R training epoch, from which we sample $K$ reasoning trajectories $o=\{o_{k}\}_{k=1}^{K}$. The second KL divergence term is approximated by
${D}_{\text{KL}}\left(\pi_{\theta}(o_{k}|c,x_{i})\|\pi_{\text{ref}}(o_{k}|c,x_{i})\right)=\frac{\pi_{\text{ref}}(o_{k}|c,x_{i})}{\pi_{\theta}(o_{k}|c,x_{i})}-\log\frac{\pi_{\text{ref}}(o_{k}|c,x_{i})}{\pi_{\theta}(o_{k}|c,x_{i})}-1,$ (6)
incorporated to ensure that the updated policy $\pi_{\theta}(\cdot|c,x_{i})$ does not deviate excessively from $\pi_{\text{ref}}(\cdot|c,x_{i})$. $\epsilon$ is the clipping threshold to prevent large and potentially destabilizing updates to the policy. The coefficient $\beta$ serves as a balancing factor between the reward-weighted likelihood term and the KL regularization term.
We conclude this section by highlighting the key strengths of our VisualQuality-R1. First, VisualQuality-R1 inherits all the advantages of ranking-based NR-IQA models, enabling effective multi-dataset training, active fine-tuning, and continual model adaptation without requiring perceptual scale realignment mikhailiuk2021consolidated , a feature notably absent in regression-based NR-IQA approaches. Second, trained via RL2R, VisualQuality-R1 mitigates the scalability and overfitting issues inherent in SFT-based models. Third, VisualQuality-R1 promises to both improve model generalizability and furnish contextually rich textual justifications alongside numerical quality scores, thereby boosting its practical relevance in real-world IQA applications.
4 Experiments
To validate VisualQuality-R1, we conduct comprehensive experiments across diverse distortion scenarios, ablation studies on key design components, and in-depth analysis of model behaviors.
Table 2: SRCC and PLCC results of NR-IQA models trained on KADID-10K. Exceptions include Q-Insight and VisualQuality-R1, which use a combined training set (KADID-10K and SPAQ). Top two results are highlighted in bold and underline, respectively.
Method Imaging-Related Distortion Processing-Related Distortion Avg
BID CLIVE KonIQ SPAQ
De-
blurring
Super-
Res.
De-
hazing
Image
Gen.
SRCC
Handcrafted
NIQE mittal2012making 0.515 0.450 0.421 0.676 0.360 0.557 0.343 0.533 0.482
BRISQUE mittal2012no 0.522 0.314 0.385 0.614 0.389 0.482 0.242 0.497 0.431
Discriminative Deep-Learning-based
UNIQUE zhang2021uncertainty 0.412 0.470 0.649 0.751 0.669 0.649 0.577 0.608 0.598
MUSIQ ke2021musiq 0.327 0.284 0.473 0.720 0.656 0.404 0.458 0.494 0.477
MANIQA yang2022maniqa 0.420 0.487 0.213 0.745 0.726 0.263 0.608 0.422 0.486
VLM-based
LIQE zhang2023blind 0.677 0.719 0.684 0.815 0.797 0.743 0.646 0.653 0.717
Q-Align wu2024qalign 0.576 0.554 0.573 0.767 0.761 0.684 0.455 0.682 0.632
DeQA-Score you2025teaching 0.702 0.743 0.677 0.852 0.785 0.710 0.643 0.738 0.731
Qwen2.5-VL-7B bai2025qwen2 0.711 0.733 0.754 0.848 0.820 0.603 0.458 0.735 0.708
Q-Insight li2025qinsight 0.784 0.761 0.806 0.872 0.831 0.724 0.601 0.749 0.766
Q-Insight 0.806 0.804 0.812 0.907 0.846 0.700 0.539 0.657 0.759
VisualQuality-R1 0.790 0.750 0.830 0.875 0.838 0.756 0.598 0.775 0.777
VisualQuality-R1 0.811 0.811 0.855 0.913 0.845 0.752 0.588 0.754 0.791
PLCC
Handcrafted
NIQE mittal2012making 0.527 0.494 0.439 0.683 0.376 0.587 0.482 0.560 0.519
BRISQUE mittal2012no 0.528 0.362 0.400 0.624 0.380 0.556 0.468 0.541 0.482
Discriminative Deep-Learning-based
UNIQUE zhang2021uncertainty 0.385 0.472 0.590 0.708 0.654 0.668 0.578 0.581 0.580
MUSIQ ke2021musiq 0.280 0.325 0.435 0.666 0.563 0.441 0.455 0.434 0.450
MANIQA yang2022maniqa 0.512 0.571 0.257 0.753 0.728 0.243 0.663 0.448 0.522
VLM-based
LIQE zhang2023blind 0.680 0.726 0.652 0.814 0.712 0.775 0.661 0.653 0.709
Q-Align wu2024qalign 0.651 0.643 0.612 0.779 0.802 0.713 0.525 0.705 0.679
DeQA-Score you2025teaching 0.743 0.795 0.703 0.858 0.838 0.763 0.688 0.790 0.772
Qwen2.5-VL-7B bai2025qwen2 0.725 0.760 0.810 0.854 0.852 0.653 0.553 0.810 0.752
Q-Insight li2025qinsight 0.796 0.795 0.829 0.872 0.857 0.798 0.669 0.810 0.803
Q-Insight 0.818 0.837 0.809 0.912 0.861 0.779 0.626 0.705 0.793
VisualQuality-R1 0.806 0.794 0.840 0.878 0.872 0.825 0.651 0.843 0.814
VisualQuality-R1 0.820 0.844 0.870 0.917 0.879 0.824 0.674 0.820 0.831
4.1 Experimental Setups
Competing Models and Training Details
Competing methods encompass three categories: 1) handcrafted models: NIQE mittal2012making and BRISQUE mittal2012no ; 2) discriminative deep-learning-based models: UNIQUE zhang2021uncertainty , MUSIQ ke2021musiq , and MANIQA yang2022maniqa ; 3) VLM-based models: LIQE zhang2023blind , Q-Align wu2024qalign , DeQA-Score you2025teaching , Q-Insight li2025qinsight , as well as the pre-trained Qwen2.5-VL-7B bai2025qwen2 baseline.
We fine-tune Qwen2.5-VL-7B bai2025qwen2 as the backbone for VisualQuality-R1 using GRPO shao2024deepseekmath . The AdamW optimizer loshchilov2017decoupled is employed with an initial learning rate of $1\times 10^{-6}$ and a linear decay schedule. For GRPO, we generate six candidate responses per prompt (i.e., $K=6$) and set the balance coefficient $\beta$ to $0.04$. Training runs on $16$ NVIDIA A100 GPUs with a minibatch size of eight per GPU, taking approximately five hours for a total of $10$ epochs.
4.2 Main Results
Single-Dataset Training
We first train NR-IQA models on the synthetic KADID-10K lin2019kadid training set ($6:2:2$ split while ensuring content independence) and test in a zero-shot setting across eight datasets with distortions arising from digital imaging and (post-)processing stages: BID ciancio2010no , CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SPAQ fang2020perceptual , Liu13 (deblurring) liu2013no , SRIQA-Bench (super-resolution) chen2025toward , Min19 (dehazing) min2019quality , and AGIQA-3K (image generation) li2023agiqa .
The SRCC and PLCC results presented in Table 2 reveal several key observations. First, all VLM-based models outperform traditional and discriminative deep-learning-based ones, with the base Qwen2.5-VL-7B achieving an SRCC of $0.708$ despite no IQA-specific training. This underscores the power of current VLMs in capturing generalizable quality cues. Second, reasoning-induced models such as Q-Insight and VisualQuality-R1 surpass SFT-based counterparts like Q-Align and DeQA-Score. Third, the proposed VisualQuality-R1 achieves the best results on average, validating that RL2R aligns better with human perception of image quality than regression-based approaches.
Multi-Dataset Training
Our RL2R approach enables multi-dataset training without the need for perceptual scale realignment. To exploit this, we train VisualQuality-R1 on a combination of KADID-10K lin2019kadid and SPAQ fang2020perceptual (again $6:2:2$ split while ensuring content independence). As shown in Table 2, VisualQuality-R1 yields consistent performance gains. Despite a minor dip in the image generation scenario, the average SRCC/PLCC rises from $0.777/0.814$ to $0.791/0.831$. In stark contrast, Q-Insight li2025qinsight fails to benefit from multi-dataset training due to its inability to address perceptual scale variations [^2] To enable multi-dataset training of Q-Insight, we linearly rescale MOSs from different IQA datasets to $[1,5]$, and apply a dataset-agnostic threshold to compute binary rewards.: KADID-10K uses ratings from $1$ to $5$, while SPAQ spans 0 to $100$.
Table 3: PLCC results of VisualQuality-R1 with varying $K$ in GRPO. The default setting is highlighted in bold.
#Generated Responses Imaging-Related Distortion Processing-Related Distortion Avg
BID CLIVE KonIQ SPAQ
De-
blurring
Super-
Res.
De-
hazing
Image
Gen.
$K=4$ 0.805 0.795 0.839 0.875 0.875 0.815 0.643 0.844 0.811
$K=5$ 0.806 0.804 0.840 0.879 0.867 0.826 0.639 0.840 0.813
$\mathbf{K=6}$ 0.806 0.794 0.840 0.878 0.872 0.825 0.651 0.843 0.814
4.3 Ablation Studies
Effect of $K$ in GRPO
We vary the number of generated responses, $K$, while keeping all other settings fixed during GRPO. Table 3 shows that reducing $K$ from six (default) to four or five has only a marginal effect, offering a favorable trade‑off between computational cost and accuracy.
Binary Reward vs. Continuous Fidelity Reward
Table 4 shows that, within the same RL2R framework, our continuous fidelity reward generalizes better than the binary reward adopted in GRPO shao2024deepseekmath . Moreover, both reward variants consistently outperform the regression‐based Q-Insight li2025qinsight , underscoring the effectiveness of our RL2R optimization.
Table 4: Comparison of different Thurstone model variants thurstone1927law in GRPO of VisualQuality-R1.
Method Imaging-Related Distortion Processing-Related Distortion Avg
BID CLIVE KonIQ SPAQ
De-
blurring
Super-
Res.
De-
hazing
Image
Gen.
SRCC
Q-Insight li2025qinsight 0.784 0.761 0.806 0.872 0.831 0.724 0.601 0.749 0.766
Binary Reward 0.780 0.756 0.821 0.877 0.834 0.748 0.587 0.771 0.772
Probability Average (Eq. (7)) 0.785 0.761 0.836 0.875 0.835 0.747 0.574 0.775 0.774
Fixed Variance of One 0.778 0.750 0.818 0.871 0.830 0.744 0.606 0.760 0.770
VisualQuality-R1 0.790 0.750 0.830 0.875 0.838 0.756 0.598 0.775 0.777
PLCC
Q-Insight li2025qinsight 0.796 0.795 0.829 0.872 0.857 0.798 0.669 0.810 0.803
Binary Reward 0.790 0.792 0.833 0.876 0.867 0.825 0.646 0.840 0.809
Probability Average (Eq. (7)) 0.796 0.797 0.844 0.875 0.861 0.817 0.621 0.831 0.805
Fixed Variance of One 0.791 0.785 0.817 0.873 0.852 0.802 0.655 0.818 0.799
VisualQuality-R1 0.806 0.794 0.840 0.878 0.872 0.825 0.651 0.843 0.814
Thurstone Model Variants
To evaluate the effectiveness of mean quality computation in Eq. (1), we compare it with an alternative that averages probabilities across individual quality comparisons:
${p}_{k}(x_{i},x_{j})=\frac{1}{K}\sum_{k^{\prime}=1}^{K}{p}_{k,k^{\prime}}(x_{i},x_{j})=\frac{1}{K}\sum_{k^{\prime}=1}^{K}\Phi\left(\frac{q_{k}(x_{i})-q_{k^{\prime}}(x_{j})}{\sqrt{\sigma^{2}(q(x_{i}))+\sigma^{2}(q(x_{j}))+\gamma}}\right).$ (7)
As reported in Table 4, averaging quality scores rather than probabilities yields higher performance across distortion scenarios, indicating more reliable comparative probability estimates and reward assignments. Taking a step further, we fix the variances in Eq. (7) to one—reducing the model to Thurstone Case V thurstone1927law . The constant-variance simplification degrades performance on nearly all datasets. This provides a strong indication that sample variances are capable of capturing the perceptual difficulty of image pairs, thus improving comparison reliability and stabilizing fidelity reward computation. Together, these findings verify that RL2R effectively embeds the Thurstone model within GRPO.
4.4 Further Analysis
Figure 3: Prediction variability decreases during GRPO. We randomly select 2020 images from each of CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SRIQA-Bench chen2025toward , and AGIQA-3K li2023agiqa . At successive training steps, we generate multiple responses per image, compute the std of the predicted quality scores, and plot the average std across images. The uniformly downward trend confirms that VisualQuality-R1 becomes steadily more stable in assessing image quality as training progresses.
Predicted Score Variability over Iterations
We randomly sample $20$ images from each of CLIVE ghadiyaram2015massive , KonIQ-10k hosu2020koniq , SRIQA-Bench (super-resolution) chen2025toward , and AGIQA-3K (image generation) li2023agiqa , respectively. At successive training checkpoints, we generate multiple responses per image and compute the standard deviation (std) of the resulting $K$ quality scores. As illustrated in Fig. 3, the std falls steadily across all datasets, indicating that predictions of VisualQuality-R1 become progressively more stable and confident.
Figure 4: Evolution of the reasoning capabilities of VisualQuality-R1 on an image super-resolved by SwinIR liang2021swinir . Initially, VisualQuality-R1 overlooks artifacts and overestimates quality; at later stages, it progressively detects stylization, blur, and filtering effects, yielding more accurate quality scores and human-aligned textual justifications. Zoom in for improved visibility.
Visual Reasoning Evolution over Iterations
Fig. 4 tracks how the visual reasoning capabilities of VisualQuality-R1 mature over the course of training. The test image is super-resolved by SwinIR liang2021swinir , which contains subtle, processing-related artifacts, making it an informative probe. Q-Insight notices that the image is “blurry” and “overexposed,” but assigns an extremely low score (i.e., $2.00$), indicating limited sensitivity to super-resolution artifacts. The base model Qwen2.5-VL-7B bai2025qwen2 swings to the opposite extreme: it praises the “clear details” and “vibrant colors,” declares the absence of blur or noise, and outputs an inflated score (i.e., $4.80$). The model clearly over-trusts superficial sharpness cues and misses hidden processing traces. In contrast, the proposed VisualQuality-R1 progressively refines its visual reasoning over iterations. At the $50$-th step, it starts to suspect artificial stylization and questions the image’s realism, yet it still values the apparent clarity. By the $200$-th step, the description becomes more balanced. It acknowledges the level of detail and clarity, yielding a slightly higher but still cautious rating. At the last step, the explanation is now decidedly nuanced. VisualQuality-R1 attributes the remaining softness to possible filtering or to the object’s inherent structure, labels the appearance “surreal,” and reduces the score to $3.00$, reflecting a judicious penalty for unnatural post-processing. In summary, RL2R guides VisualQuality-R1 from naïve, superficial remarks to sophisticated, human-aligned reasoning that correctly identifies subtle super-resolution artifacts and calibrates quality scores accordingly.
5 Conclusion and Discussion
We have introduced VisualQuality-R1, a reasoning-induced NR-IQA model optimized via RL2R. Our approach is grounded in the intrinsic relativity of visual quality, seamlessly integrating the Thurstone model within GRPO to capture predictive uncertainty. By introducing the continuous fidelity reward, VisualQuality-R1 delivers more precise policy-gradient signals.
Extensive experiments validate that VisualQuality-R1 consistently surpasses strong discriminative deep learning-based methods and a reasoning-induced baseline. Notably, it bridges the performance gap between synthetic and realistic distortions, demonstrating robustness to dataset inductive biases and noise. In addition to quantitative improvements, VisualQuality-R1 generates contextually rich, human-aligned quality descriptions, which not only enhance transparency and interpretability but also facilitate user trust and post-hoc diagnosis in downstream tasks, such as content filtering, local enhancement prioritization, and quality-aware image retrieval.
Limitations and Future Directions
Despite the generalization capabilities demonstrated by VisualQuality-R1, several limitations and promising research directions merit further discussion. First, as a specific case of test-time scaling, VisualQuality-R1 is slow, expensive, and memory-hungry; it may also compound early errors into confidently wrong predictions. It is thus desirable to incorporate sample-adaptive reasoning, rationale compression or distillation, and self-consistency sampling to make VisualQuality-R1 faster, cheaper, and more robust. Second, VisualQuality-R1 relies on a single, fixed text prompt for all images, regardless of the underlying distortion scenario or application context. Incorporating application-aware prompt adaptation, for example via learned prompt-tuning or dynamic template selection, could tailor VisualQuality-R1’s reasoning and scoring to specific image processing tasks, therefore improving its flexibility and accuracy. Third, VisualQuality-R1 is currently formulated as an NR-IQA model, focusing solely on distorted inputs without access to pristine-quality counterparts. It is interesting to extend VisualQuality-R1 to a reference-based setting, which allows a (possibly corrupted) reference image wu2024comprehensive —potentially differing in resolution, color gamut, dynamic range, or bit depth—to serve as a flexible anchor for content fidelity. Last, we foresee adapting the proposed RL2R learning algorithm to other perceptual assessment tasks, including image aesthetics assessment tang2013content , human age estimation guo2009human , and perceptual similarity ranking fu2023dreamsim . Collectively, these promising directions aspire to foster more intelligent, transparent, and adaptable perceptual systems.
Acknowledgments
This work was supported in part by the Hong Kong ITC Innovation and Technology Fund (9440379 and 9440390) and the PolyU-OPPO Joint Innovative Research Center.
Appendix
Appendix A Generalization Probing via gMAD Competition
The group maximum differentiation (gMAD) competition ma2018group is a model comparison framework designed to evaluate the generalization capability of computational models, particularly in scenarios where exhaustive ground-truth labeling is impractical. In this framework, models take turns serving as the defender, while the remaining models act as attackers. The attackers aim to identify image pairs that the defender assigns similar quality scores to, but which they themselves rate very differently. These “adversarial” pairs are subsequently assessed by human observers, allowing for an efficient and targeted examination of model strengths and weaknesses.
Applying gMAD to compare KADID-10K-trained VisualQuality-R1 and Q-Insight on KonIQ-10k hosu2020koniq reveals a clear performance distinction. VisualQuality-R1 consistently uncovers perceptual inconsistencies in Q-Insight’s predictions. Conversely, when subjected to attacks, VisualQuality-R1 demonstrates strong robustness, aligning closely with human perception of image quality. This dual capability—high aggressiveness in exposing flaws and strong resistance to adversarial challenges—underscores VisualQuality-R1’s superior generalization, affirming its reliability for real-world IQA.
Appendix B Model Complexity Comparison
We present a quantitative comparison of model complexity across representative NR-IQA methods. As shown in Table 5, VisualQuality-R1 incurs higher inference costs than discriminative models such as MUSIQ ke2021musiq and MANIQA yang2022maniqa , as well as small-scale VLMs like LIQE zhang2023blind . This increased computational demand, however, reflects a deliberate design choice to support fine-grained quality reasoning and robust quality rating across diverse distortion scenarios.
Figure 5: gMAD competition results between VisualQuality-R1 and Q-Insight li2025qinsight . (a) Fixed Q-Insight at the low-quality level. (b) Fixed Q-Insight at the high-quality level. (c) Fixed VisualQuality-R1 at the low-quality level. (d) Fixed VisualQuality-R1 at the high-quality level.
Table 5: Model complexity comparison using a $512\times 384\times 3$ image as input.
Method #Parameters
Inference
Time
Inference
Memory
#TFLOPs
UNIQUE zhang2021uncertainty 22.32 M 0.02 s 1.62 G 0.029
MUSIQ ke2021musiq 27.13 M 0.05 s 1.69 G 0.026
MANIQA yang2022maniqa 135.75 M 0.03 s 2.13 G 0.217
LIQE zhang2023blind 151.28 M 0.03 s 2.15 G 0.131
Q-Align wu2024qalign 8.20 B 0.14 s 17.1 G 1.98
DeQA-Score you2025teaching 8.20 B 0.11 s 17.1 G 1.98
Q-Insight li2025qinsight 8.29 B 2.72 s 17.6 G 8.71
VisualQuality-R1 8.29 B 2.34 s 17.6 G 7.74
Figure 6: Visual reasoning samples from VisualQuality-R1 under the synthetic distortion scenario.
Figure 7: Visual reasoning samples from VisualQuality-R1 under the realistic distortion scenario.
Appendix C Additional Visual Reasoning Examples
We provide additional visual reasoning examples to illustrate the perceptual alignment of VisualQuality-R1 under both synthetic and realistic distortion scenarios. It should be noted that VisualQuality-R1 occasionally assigns quality scores outside the prescribed range of $\left[1,5\right]$ (e.g., a score of $0.2$ for the second example in Fig. 6), which reflects a common limitation in LLMs. Nevertheless, these deviations have negligible impact on the overall quality prediction performance.